# jXBW: Fast Substructure Search in Large-Scale JSONL Datasets for Foundation Model Applications

### v1

@@ Abstract.
XBW Transform, Substructure Search, JSONL, Compressed Data Structure, Foundation Models
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Data structuresccs: Information systems Query processingccs: Theory of computation Data compressionccs: Computing methodologies Machine learning
@@ 1. Introduction
JSON Lines (JSONL) is a widely adopted data format in which each line contains a single, self-contained JSON object. The importance of JSONL has grown significantly with the rise of Foundation Models (Bommasani et al., 2021), where structured prompt inputs containing diverse, semi-structured information must be efficiently processed and queried. Foundation Models such as GPT (Brown et al., 2020; Radford et al., 2019), BERT (Devlin et al., 2018), T5 (Raffel et al., 2020), LLaMA (Touvron et al., 2023a, b), and Claude (Anthropic, 2023) are commonly deployed with JSONL-formatted prompt inputs for tasks including question answering, text generation, and information retrieval, with recent empirical studies demonstrating up to 40 % performance improvements over plain text formats (Chen et al., 2024a) and comparative studies with GPT-4o showing that JSON-formatted inputs achieve the highest accuracy for complex structured data processing compared to YAML or hybrid CSV formats (Zhang et al., 2025). Enterprise applications and research platforms commonly employ JSONL format to manage complex prompt datasets that can contain millions of structured queries and contextual information (Chen et al., 2024b; Liu et al., 2021; Wang et al., 2022). There is therefore a compelling and growing need to develop novel, more powerful methods for effectively utilizing large-scale JSONL datasets to draw out the capabilities of foundation models through prompt engineering.
Substructure search in JSONL data is essential for prompt engineering in foundation models. The goal is to find all JSON objects that contain a given query as a substructure. Each JSON object can be naturally represented as a tree structure, transforming substructure search into a tree matching problem. This problem must handle both ordered structures (JSON arrays) and unordered structures (JSON objects). Simple pointer-based approaches require traversing each JSON tree individually and performing explicit enumeration and comparison of child nodes, making them computationally prohibitive for large-scale datasets. While space-efficient succinct tree representations called SJSON for JSON objects exist (Lee et al., 2020), they still suffer from the same computational bottlenecks in substructure search tasks.
Alternatively, one can transform JSON objects into XML and leverage XQuery for substructure search, benefiting from mature XML indexing technologies (Raut and Atique, 2014; Bruno et al., 2002). However, this approach introduces substantial overhead: the JSON-to-XML conversion process increases data volume and processing latency, while semantic differences between JSON and XML models—particularly in handling arrays, null values, and type distinctions—complicate query formulation and result interpretation. Moreover, this approach requires performing individual substructure searches across numerous XML documents converted from each JSON object, leading to computational overhead that scales linearly with dataset size. Additionally, XML indexing techniques cannot exploit the structural similarities commonly found among JSON objects in JSONL datasets, missing optimization opportunities that are naturally available in JSON-native approaches. Therefore, an important open challenge is to develop specialized data structures and algorithms that can efficiently support substructure search directly on JSONL data.
Contribution. We present jXBW (JSONL eXtended Burrows-Wheeler Transform), a novel approach for fast substructure search on large JSONL datasets that addresses the computational bottlenecks of existing methods through three key innovations:
(1) Merged Tree Representation: We convert each JSON object into a tree structure and merge multiple trees into a single unified merged tree. This representation exploits structural similarities among JSON objects to reduce memory overhead and enables efficient substructure search.
(2) Compressed Tree Index (jXBW): Since pointer-based representations require expensive tree traversal for substructure matching, we introduce jXBW—a compressed representation based on the eXtended Burrows-Wheeler Transform (XBW) for trees (Ferragina et al., 2007). XBW enables efficient tree navigation operations in $O(\log\sigma)$ time, where $\sigma$ is the alphabet size of distinct node labels in the tree, and supports subpath search queries in $O(|P|\log\sigma)$ time for paths of length $|P|$. Crucially, subpath search can locate matching paths starting from any node, not just the root. This capability enables an efficient substructure search algorithm that avoids exhaustive tree traversal and explicit comparison of child nodes.
(3) Efficient Substructure Search Algorithm on jXBW: We present a fast substructure search algorithm specifically designed for the jXBW representation that operates in three key steps: (i) path decomposition that transforms query trees into root-to-leaf paths and uses jXBW’s subpath search to directly locate matching positions, (ii) ancestor computation that identifies common subtree root positions reachable by all query paths through efficient parent operations, and (iii) tree identifier collection that employs an adaptive strategy to automatically select optimal processing approaches based on query characteristics. This three-step approach ensures correctness while avoiding exhaustive tree traversal, where tree identifiers corresponding to JSONL line numbers are computed through precise operations.
@@ 2.1. JSONL and its Tree Representation
  • <<object>> ::= { <<members>>? }
  • <<members>> ::= <<pair>> (, <<pair>>)*
  • <<pair>> ::= <<string>> : <<value>>
Here, <<value>> may be one of the following: <<string>>, <<number>>, <<object>>, <<array>>, true, false, or null.
An <<object>> may optionally include <<members>>, which consist of one or more unordered <<pair>> elements. Each <<pair>> represents a key-value association, where the key is a <<string>>, and the value is any valid JSON data type as defined above. An <<array>> represents an ordered sequence of elements, each of which may be of any valid JSON data type.
@@
$O_{1}$ $=\{\text{"person"}:\{\text{"name"}:\text{"Alice"},\text{"age"}:30\},$
$\quad\quad\text{"hobbies"}:[\text{"reading"},\text{"cycling"}]\}$
$O_{2}$ $=\{\text{"person"}:\{\text{"name"}:\text{"Bob"},\text{"age"}:30\},$
$\quad\quad\text{"hobbies"}:[\text{"reading"}]\}$
Figure 1. Illustration of tree structures $T_{1}$ (left) and $T_{2}$ (right), derived from the example JSON objects $O_{1}$ and $O_{2}$, respectively.
Figure 2. Illustration of the merged tree $T_{M}$ derived from $T_{1}$ and $T_{2}$.
Tree Representation. Each object $O_{i}$ can be systematically transformed into a tree-based representation with node labels. Let $T_{i}$ denote the tree constructed from the $i$-th object $O_{i}$ and $|T_{i}|$ denote the number of nodes in $T_{i}$. The root node of $T_{i}$ is labeled "object" and has as many unordered child nodes as there are <<pair>> elements in the <<members>>. If the <<object>> contains no <<members>>, the root node has no children.
Each <<pair>> consists of a key-value association of the form <<string>>:<<value>>. In the corresponding tree structure, each child of the root is labeled with the key <<string>> and has a single child node representing the value. The value node may correspond to any valid JSON type: <<string>>, <<number>>, <<object>>, <<array>>, true, false, or null. If the value is an <<object>>, the tree expands recursively, treating the value node as the root of a new subtree. If the value is an <<array>>, the value node has multiple children, each representing an element of the array in the specified order. Each element may itself be of any valid JSON type.
Furthermore, each leaf node $\ell$ in tree $T_{i}$ is annotated with the tree identifier $i$ to track its origin. For any leaf node $\ell$, we define $ids(\ell)$ as the set of all tree identifiers associated with that leaf node. In individual trees, this is simply $ids(\ell)=\{i\}$ for leaf node $\ell$ in tree $T_{i}$.
Fig. 1 illustrates the tree structures $T_{1}$ and $T_{2}$ derived from the example JSONL objects $O_{1}$ and $O_{2}$ which describe two individuals: Alice and Bob. The leaf nodes in $T_{1}$ and $T_{2}$ are annotated with identifiers $1$ and $2$, respectively.
@@
  • Primitive values: exact value comparison
A naive approach to substructure search in problem 2.1 would involve searching each individual tree $T_{i}$ separately for the query tree $Q$. However, this approach has significant computational drawbacks. First, it requires $O(N\times\sum_{i=1}^{N}|T_{i}|\times|Q|)$ time complexity, where $|T_{i}|$ is the number of nodes in $T_{i}$ and each tree must be traversed independently to find potential matches. Second, this approach fails to exploit structural similarities between trees, resulting in redundant computations when many trees share common substructures. To overcome these limitations, we introduce an efficient merged tree representation that consolidates shared structural patterns while preserving individual tree identities.
@@ 3. Merged Tree Representation
An example of the merged tree $MT$ derived from $T_{1}$ and $T_{2}$ is presented in Fig. 2. In this example, both the leaf nodes labeled “30” and “reading” are reached through paths shared by $T_{1}$ and $T_{2}$, resulting in these leaf nodes $\ell_{3}$ and $\ell_{4}$ having $ids(\ell_{3})=\{1,2\}$ and $ids(\ell_{4})=\{1,2\}$ in the merged tree. The other leaf nodes $\ell_{1}$ labeled “Alice”, $\ell_{2}$ labeled “Bob”, and $\ell_{5}$ labeled “cycling” have $ids(\ell_{1})=\{1\}$, $ids(\ell_{2})=\{2\}$, and $ids(\ell_{5})=\{1\}$, respectively.
@@ 3.1. Substructure Search on Merged Tree
For the example query object $\{\text{"name"}:\text{"Bob"},\text{"age"}:30\}$, we convert this object into the tree structure $Q$ as shown in Fig. 3. The algorithm searches for nodes labeled ”Object” by recursively traversing $MT$ in Fig. 2 during the candidate finding phase. It finds root and internal nodes labeled ”Object” in $MT$, and recursively checks whether the subtree rooted at each node matches the structure of $Q$ during the substructure matching phase. One subtree rooted at an internal node labeled ”Object” is found to match the structure of $Q$. The sets of identifiers $I(\ell_{2})=\{2\}$ and $I(\ell_{3})=\{1,2\}$ associated with the leaves $\ell_{2}$ and $\ell_{3}$ in $MT$ are collected. In the solution identification phase, we compute the intersection of all collected identifier sets as $I(\ell_{2})\cap I(\ell_{3})=\{2\}\cap\{1,2\}=\{2\}$. Thus the solution IDs are $\{2\}$.
Complexity. The substructure search algorithm on the merged tree has significant efficiency limitations. The time complexity is $O(|MT|\times|Q|)$ in the worst case. This occurs when the algorithm must examine every node in $MT$ as a potential starting point for matching the root of $Q$, and for each candidate node, recursively compare the entire structure of $Q$ against the corresponding subtree in $MT$. Furthermore, during the substructure matching phase, for each pair of nodes (one from the substructure in $MT$ and one from $Q$), the algorithm must exhaustively search through all children of the $MT$ node to find those with matching labels to the children of the $Q$ node. Although this approach is more efficient than individual searches on each of the $N$ original trees (which would require $O(N\times\sum_{i=1}^{N}|T_{i}|\times|Q|)$ time, the need to traverse the entire merged tree remains a computational bottleneck, especially for large JSONL datasets with complex tree structures. To overcome these efficiency limitations and enable fast substructure queries on large-scale data, we present an efficient substructure search algorithm based on jXBW in the following sections.
@@ 4.1. Wavelet Tree and Matrix
Figure 4. Illustration of the construction process of the XBW representation of the merged tree $MT^{\prime}$. Note array $A_{anc}$ is not included in the final jXBW representation and is displayed for clarity.
@@
The construction of the XBW representation for our merged tree $MT$ follows a systematic process as illustrated in Fig. 4.
@@
Step 1: Symbol Table Creation and Label Normalization. Since $MT$ contains diverse node labels from the original JSONL objects (e.g., field names like “name”, “age”, string values like “Alice”, “Bob”), we first establish a symbol table that creates a bijective mapping from each unique node label to a symbol in a unified alphabet $\Sigma$ (Fig. 4(I)). This mapping enables efficient lexicographic comparison by replacing potentially long string labels with compact symbols.
@@
Using this mapping, all node labels in $MT$ are converted to the standardized alphabet $\Sigma$ (Fig. 4(II)). Let $MT^{\prime}$ denote the merged tree with normalized labels and lexicographically sorted children.
@@
Step 2: Array Construction via DFS Traversal. The XBW representation consists of five synchronized arrays (collectively referred to as jXBW arrays), each of length $|MT|$, constructed via a depth-first traversal of $MT^{\prime}$ (Fig. 4(III)):
  • $A_{label}[i]$: stores the label of the $i$-th visited node during DFS traversal.
@@
  • $A_{anc}[i]$: stores the sequence of node labels on the path from the root to the parent of the $i$-th visited node. For the root node, $A_{anc}[1]=\varepsilon$ (empty string).
@@
  • $A_{last}[i]$: a binary array where $A_{last}[i]=1$ if the $i$-th visited node is the rightmost child of its parent, and 0 otherwise. For the root, $A_{last}[1]=1$.
@@
  • $A_{leaf}[i]$: a binary array where $A_{leaf}[i]=1$ if the $i$-th visited node is a leaf, and 0 otherwise.
@@
  • $A_{ids}[i]$: stores the set of identifiers idsids associated with the $i$-th visited node.
@@
Step 3: Lexicographic Sorting. All five arrays $A_{label}$, $A_{anc}$, $A_{last}$, $A_{leaf}$, and $A_{ids}$ are simultaneously sorted based on the lexicographic order of the sequences in $A_{anc}$ (Fig. 4(IV)). In addition, the array $A_{label}$ is indexed by the wavelet matrix. The binary arrays $A_{last}$ and $A_{leaf}$ are also indexed by the rank and select dictionary.
@@
To reduce memory usage, $A_{ids}$ is compacted to contain only entries corresponding to leaf nodes, arranged consecutively without gaps. For a leaf node at array position $i$, its tree identifiers are accessed via $A_{ids}[\text{rank}_{1}(A_{leaf},i)]$, where the rank operation maps the leaf position to the corresponding index in the compacted array.
@@
The resulting arrays, excluding $A_{anc}$, constitute the XBW representation of $MT$, which provides a compressed representation that maintains the essential structural information enabling tree navigation operations and a subpath search query while reducing space requirements compared to the explicit tree representation. Note that array $A_{anc}$ is not included in the final XBW representation, but is used only during the construction process. Thus, the space complexity of jXBW is $O(|MT^{\prime}|\log{\sigma})$. While this may be comparable to the pointer-based representation $O(|MT^{\prime}|\log|MT^{\prime}|)$ in terms of space complexity for JSONL datasets with large vocabularies, jXBW provides significant advantages in query performance.
@@
An important property of the jXBW representation is that (i) siblings are contiguously stored in arrays in lexicographic order of their labels; (ii) the element in $A_{last}$ corresponding to the rightmost child is always $1$. These properties are ensured by the XBW construction process: (a) the depth-first traversal of $MT^{\prime}$; (b) the stable lexicographic sorting based on $A_{anc}$; (c) the binary encoding in $A_{last}$, where each sibling block is encoded as a sequence of 0s terminated by a $1$ for the rightmost child.
@@
For example, in Fig. 4, the contiguous ranges of array positions of siblings for nodes labeled ’A’ (root), ’A’ (internal), ’B’, ’C’, ’F’, ’G’, ’I’ are [2,3], [4,5], [6], [7,8], [9], [10], [11,12], respectively. Array positions 3, 5, 6, 8, 9, 10, 12, which correspond to the rightmost child of each node, are marked as $1$ in $A_{last}$.

### v2

# jXBW: Fast Substructure Search for Large-Scale JSONL Datasets with LLM Applications

@@ Abstract.
XBW Transform, Substructure Search, JSONL, Compressed Data Structure, Large Language Model (LLM), Retrieval-Augmented Generation (RAG)
copyright: acmlicensedjournalyear: 2026doi: XXXXXXX.XXXXXXXconference: 2026 International Conference on Management of Data; May 31–June 5, 2026; Bengaluru, Indiabooktitle: 2026 International Conference on Management of Dataisbn: 978-1-4503-XXXX-X/2026/05ccs: Information systems Data structuresccs: Information systems Query processingccs: Theory of computation Data compressionccs: Computing methodologies Machine learning
@@ 1. Introduction
Recent studies also underscore the role of prompt formats in LLM performance. He et al. (He et al., 2024) show that the choice of format — JSON versus Markdown — can boost accuracy by as much as 40% in reasoning tasks. Similarly, Zhang et al. (Zhang et al., 2025) demonstrate that structured JSON prompts outperform YAML or hybrid CSV in GPT-4o, yielding the most reliable results for complex tasks. These findings make clear that structured inputs are not only more efficient but also more accurate, strengthening the case for JSONL as a unifying format for modern AI and data systems.
In enterprise applications, data are often stored across multiple relational databases in normalized form, requiring complex joins to assemble the information needed by downstream applications. While relational databases excel at ensuring consistency, such normalized schemas are not well aligned with language models, which benefit from flattened, object-like representations. A common approach is therefore to transform relational data into JSON objects, enabling easier integration with LLMs. This reflects a broader trend: object databases and modern NoSQL systems (e.g., MongoDB) store data natively as JSON, offering a natural fit for LLM-based analytics (Liu et al., 2021; Wang et al., 2022). These considerations further highlight why JSONL has emerged as a practical and widely adopted format for large-scale AI and data science workflows.
Despite its adoption, efficient querying over JSONL remains a critical bottleneck. In many applications, the task is not simply to retrieve documents by keyword or field but to identify all objects that contain a given substructure (i.e., substructure search). Examples include: Drug discovery, where chemists must query millions of compounds for molecules containing specific functional groups; Prompt engineering, where developers need to extract all prompts containing a particular schema fragment or contextual pattern; Geospatial analytics, where analysts search for OSM entities with nested attributes (e.g., convenience stores with parking lots).
Such substructure search queries are computationally expensive. Each JSON object can be naturally represented as a tree, making substructure search essentially a tree-matching problem. Direct traversal therefore requires exhaustive matching and scales linearly with dataset size. The problem is further complicated by the need to handle both ordered structures (JSON arrays) and unordered structures (JSON objects). Succinct JSON representations such as SJSON (Lee et al., 2020) reduce memory usage but provide no acceleration for substructure queries. Alternatively, one can transform JSON objects into XML and leverage XQuery for substructure search, benefiting from mature XML indexing technologies (Raut and Atique, 2014; Bruno et al., 2002). However, this approach introduces substantial overhead: the JSON-to-XML conversion process increases data volume and processing latency, while semantic differences between JSON and XML models—particularly in handling arrays, null values, and type distinctions—complicate query formulation and result interpretation. Moreover, this approach requires performing individual substructure searches across numerous XML documents converted from each JSON object, leading to computational overhead that scales linearly with dataset size.
The challenge is particularly pressing in the context of retrieval-augmented generation (RAG). RAG has become a central paradigm for improving factuality in LLMs by injecting retrieved knowledge into the generation process (Lewis et al., 2020). Yet conventional RAG methods operate on unstructured text chunks, which hinders multi-hop reasoning and global integration. To overcome this, recent work has emphasized structurization: Retrieval-And-Structuring (RAS) (Jiang et al., 2025) dynamically builds query-specific knowledge graphs, while StructRAG (Li et al., 2024) selects task-appropriate structures such as tables or graphs. These advances share a key insight: structure is essential for robust reasoning. However, existing methods focus on natural language documents. For semi-structured data such as JSONL, no efficient system supports structure-aware retrieval and reasoning. Therefore, an important open challenge is to develop specialized data structures and algorithms that can efficiently support substructure search directly on JSONL data.
Our contributions are as follows:
  • We identify the JSONL substructure search problem and show why existing JSON/XML indexing approaches fail to address it efficiently.
  • We propose jXBW, a novel compressed index combining (i) a merged tree representation, (ii) an XBW-based succinct tree index, and (iii) a three-phase substructure search algorithm.
  • We provide a complexity analysis, showing that jXBW achieves query-dependent complexity, determined by query structure rather than dataset size, while keeping space usage succinct.
[…]
This establishes jXBW as the first practical solution for JSONL substructure search, opening up new opportunities for large-scale LLM-based analytics.
@@ 2.1. JSONL and its Tree Representation
<object> ::= { <members> ? }
<members> ::= <pair> ( , <pair> )*
<pair> ::= <string> : <value>
Here, <value> may be one of the following: <string>, <number>, <object>, <array>, true, false, or null.
An <object> may optionally include <members>, which consist of one or more unordered <pair> elements. Each <pair> represents a key-value association, where the key is a <string>, and the value is any valid JSON data type as defined above. An <array> represents an ordered sequence of elements, each of which may be of any valid JSON data type.
@@
O1 = {"person":{"name":"Alice","age":30}, "hobbies":["reading","cycling"]}
O2 = {"person":{"name":"Bob","age":30}, "hobbies":["reading"]}
Figure 1. Tree structures $T_{1}$ (left) and $T_{2}$ (right), derived from the example JSON objects $O_{1}$ and $O_{2}$, respectively.
Figure 2. Merged tree $T_{M}$ derived from $T_{1}$ and $T_{2}$.
Tree Representation. Each object $O_{i}$ can be systematically transformed into a tree-based representation with node labels. Let $T_{i}$ denote the tree constructed from the $i$-th object $O_{i}$ and $|T_{i}|$ denote the number of nodes in $T_{i}$. The root node of $T_{i}$ is labeled “object” and has as many unordered child nodes as there are <pair> elements in the <members>. If the <object> contains no <members>, the root node has no children.
Each <pair> consists of a key-value association of the form <string>:<value>. In the corresponding tree structure, each child of the root is labeled with the key <string> and has a single child node representing the value. The value node may correspond to any valid JSON type: <string>, <number>, <object>, <array>, true, false, or null. If the value is an <object>, the tree expands recursively, treating the value node as the root of a new subtree. If the value is an <array>, the value node has multiple children, each representing an element of the array in the specified order. Each element may itself be of any valid JSON type.
Furthermore, each leaf node $\ell$ in tree $T_{i}$ is annotated with the tree identifier $i$ to track its origin. For any leaf node $\ell$, we define $\mathrm{ids}(\ell)$ as the set of all tree identifiers associated with that leaf node. In individual trees, this is simply $\mathrm{ids}(\ell)=\{i\}$ for leaf node $\ell$ in tree $T_{i}$.
Figure 1 illustrates the tree structures $T_{1}$ and $T_{2}$ derived from the example JSONL objects $O_{1}$ and $O_{2}$ which describe two individuals: Alice and Bob. The leaf nodes in $T_{1}$ and $T_{2}$ are annotated with identifiers $1$ and $2$, respectively.
@@
  • Primitive values: exact match under JSON typing (strings, numbers, booleans, and null)
A straightforward approach to substructure search in problem 2.1 would involve searching each individual tree $T_{i}$ separately for the query tree $Q$. However, this approach has significant computational drawbacks. First, it requires $O\!\big((\sum_{i=1}^{N}|T_{i}|)\cdot|Q|\big)$ time complexity, where $|T_{i}|$ is the number of nodes in $T_{i}$ and each tree must be traversed independently to find potential matches. Second, this approach fails to exploit structural similarities between trees, resulting in redundant computations when many trees share common substructures. To overcome these limitations, we introduce an efficient merged tree representation that consolidates shared structural patterns while preserving individual tree identities.
@@ 3. Merged Tree Representation
An example of the merged tree $MT$ derived from $T_{1}$ and $T_{2}$ is presented in Figure 2. In this example, both the leaf nodes labeled “30” and “reading” are reached through paths shared by $T_{1}$ and $T_{2}$, resulting in these leaf nodes $\ell_{3}$ and $\ell_{4}$ having $ids(\ell_{3})=\{1,2\}$ and $ids(\ell_{4})=\{1,2\}$ in the merged tree. The other leaf nodes $\ell_{1}$ labeled “Alice”, $\ell_{2}$ labeled “Bob”, and $\ell_{5}$ labeled “cycling” have $ids(\ell_{1})=\{1\}$, $ids(\ell_{2})=\{2\}$, and $ids(\ell_{5})=\{1\}$, respectively.
@@ 3.1. Substructure Search on Merged Tree
For the example query object $\{\text{"name"}:\text{"Bob"},\text{"age"}:30\}$, we convert this object into the tree structure $Q$ as shown in Figure 3. The algorithm searches for nodes labeled “object” by recursively traversing $MT$ in Figure 2 during the candidate finding phase. It finds root and internal nodes labeled “object” in $MT$, and recursively checks whether the subtree rooted at each node matches the structure of $Q$ during the substructure matching phase. One subtree rooted at an internal node labeled “object” is found to match the structure of $Q$. The sets of identifiers $I(\ell_{2})=\{2\}$ and $I(\ell_{3})=\{1,2\}$ associated with the leaves $\ell_{2}$ and $\ell_{3}$ in $MT$ are collected. In the solution identification phase, we compute the intersection of all collected identifier sets as $I(\ell_{2})\cap I(\ell_{3})=\{2\}\cap\{1,2\}=\{2\}$. Thus the solution IDs are $\{2\}$.
Complexity. The substructure search algorithm on the merged tree has significant efficiency limitations. The time complexity is $O(|MT|\times|Q|)$ in the worst case. This occurs when the algorithm must examine every node in $MT$ as a potential starting point for matching the root of $Q$, and for each candidate node, recursively compare the entire structure of $Q$ against the corresponding subtree in $MT$. Furthermore, during the substructure matching phase, for each pair of nodes (one from the substructure in $MT$ and one from $Q$), the algorithm must exhaustively search through all children of the $MT$ node to find those with matching labels to the children of the $Q$ node. Although this approach is more efficient than individual searches on each of the $N$ original trees (which would require $O\!\big((\sum_{i=1}^{N}|T_{i}|)\cdot|Q|\big)$ time, the need to traverse the entire merged tree remains a computational bottleneck, especially for large JSONL datasets with complex tree structures. In practice, shared prefixes substantially shrink the effective search space compared to per-tree scans, as common paths are traversed once in $MT$. To overcome these efficiency limitations and enable fast substructure queries on large-scale data, we present an efficient substructure search algorithm based on jXBW in the following sections.
@@ 4.1. Wavelet Tree and Matrix
Figure 4. Construction process of the XBW representation of the merged tree $MT^{\prime}$. Note array $A_{\mathit{anc}}$ is not included in the final jXBW representation and is displayed for clarity.
@@
The construction of the XBW representation for our merged tree $MT$ follows a systematic process as illustrated in Figure 4.
@@
Step 1: Symbol Table Creation and Label Normalization. Since $MT$ contains diverse node labels from the original JSONL objects (e.g., field names like “name”, “age”, string values like “Alice”, “Bob”), we first establish a symbol table that creates a bijective mapping from each unique node label to a symbol in a unified alphabet $\Sigma$ (Figure 4(I)). This mapping enables efficient lexicographic comparison by replacing potentially long string labels with compact symbols.
@@
Using this mapping, all node labels in $MT$ are converted to the standardized alphabet $\Sigma$ (Figure 4(II)). Let $MT^{\prime}$ denote the merged tree with normalized labels and lexicographically sorted children.
@@
Step 2: Array Construction via DFS Traversal. The XBW representation consists of five synchronized arrays (collectively referred to as jXBW arrays), each of length $|MT|$, constructed via a depth-first traversal of $MT^{\prime}$ (Figure 4(III)):
  • $A_{\mathit{label}}[i]$: stores the label of the $i$-th visited node during DFS traversal.
@@
  • $A_{\mathit{anc}}[i]$: stores the sequence of node labels on the path from the root to the parent of the $i$-th visited node. For the root node, $A_{\mathit{anc}}[1]=\varepsilon$ (empty string).
@@
  • $A_{\mathit{last}}[i]$: a binary array where $A_{\mathit{last}}[i]=1$ if the $i$-th visited node is the rightmost child of its parent, and 0 otherwise. For the root, $A_{\mathit{last}}[1]=1$.
@@
  • $A_{\mathit{leaf}}[i]$: a binary array where $A_{\mathit{leaf}}[i]=1$ if the $i$-th visited node is a leaf, and 0 otherwise.
@@
  • $A_{\mathit{ids}}[i]$: stores the set of identifiers $\mathrm{ids}$ associated with the $i$-th visited node.
@@
Step 3: Lexicographic Sorting. All five arrays $A_{\mathit{label}}$, $A_{\mathit{anc}}$, $A_{\mathit{last}}$, $A_{\mathit{leaf}}$, and $A_{\mathit{ids}}$ are simultaneously sorted based on the lexicographic order of the sequences in $A_{\mathit{anc}}$ (Figure 4(IV)). In addition, the array $A_{\mathit{label}}$ is indexed by the wavelet matrix. The binary arrays $A_{\mathit{last}}$ and $A_{\mathit{leaf}}$ are also indexed by the rank and select dictionary.
@@
To reduce memory usage, $A_{\mathit{ids}}$ is compacted to contain only entries corresponding to leaf nodes, arranged consecutively without gaps. For a leaf node at array position $i$, its tree identifiers are accessed via $A_{\mathit{ids}}[\text{rank}_{1}(A_{\mathit{leaf}},i)]$, where the rank operation maps the leaf position to the corresponding index in the compacted array.
@@
The resulting arrays, excluding $A_{\mathit{anc}}$, constitute the XBW representation of $MT$, which provides a compressed representation that maintains the essential structural information enabling tree navigation operations and a subpath search query while reducing space requirements compared to the explicit tree representation. Note that array $A_{\mathit{anc}}$ is not included in the final XBW representation, but is used only during the construction process. Thus, the space complexity of jXBW is $O(|MT^{\prime}|\log{\sigma})$. While this may be comparable to the pointer-based representation $O(|MT^{\prime}|\log|MT^{\prime}|)$ in terms of space complexity for JSONL datasets with large vocabularies, jXBW provides significant advantages in query performance.
@@
An important property of the jXBW representation is that (i) siblings are contiguously stored in arrays in lexicographic order of their labels; (ii) the element in $A_{\mathit{last}}$ corresponding to the rightmost child is always $1$. These properties are ensured by the XBW construction process: (a) the depth-first traversal of $MT^{\prime}$; (b) the stable lexicographic sorting based on $A_{\mathit{anc}}$; (c) the binary encoding in $A_{\mathit{last}}$, where each sibling block is encoded as a sequence of 0s terminated by a $1$ for the rightmost child.
@@
For example, in Figure 4, the contiguous ranges of array positions of siblings for nodes labeled ’A’ (root), ’A’ (internal), ’B’, ’C’, ’F’, ’G’, ’I’ are [2,3], [4,5], [6], [7,8], [9], [10], [11,12], respectively. Array positions 3, 5, 6, 8, 9, 10, 12, which correspond to the rightmost child of each node, are marked as $1$ in $A_{\mathit{last}}$.